# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ E.1 Gaussian consistency loss
$KL(P||Q)=\frac{1}{2}\left[(\mu_{Q}-\mu_{P})^{\top}\Sigma_{Q}^{-1}(\mu_{Q}-\mu_{P})+\mbox{tr}(\Sigma_{Q}^{-1}\Sigma_{P})-\log\frac{|\Sigma_{P}|}{|\Sigma_{Q}|}-n\right]\,.$
@@
$\mathcal{L}_{cons}=\mathbb{E}_{\mathbf{i}\sim p(\mathbf{i})}\left[KL_{z}(\hat{P}^{(\mathbf{i})}(z)|\hat{P^{(\omega(\mathbf{i}))}}(z))\right]+\mathbb{E}_{z\sim\hat{P}^{(\mathbf{i})}\left(\mathbf{Z}\right)}\left[KL_{Y}\left(\hat{P}^{(\mathbf{i})}\left({Y}|{\mathbf{Z}}=z\right)||{P}^{(0)}\left({Y}|{\mathbf{Z}}=z\right)\right)\right]\\ =\frac{1}{2}\mathbb{E}_{\mathbf{i}\sim p(\mathbf{i})}\left[\sum_{k}\left(\frac{({\mu}_{Z,k}+{\bm{\omega}}_{k}^{\top}\mathbf{i}-\hat{\mu}^{(\mathbf{i})}_{Z,k})^{2}}{\sigma_{Z,k}^{2}}+\frac{{\widehat{\sigma^{2}}}^{(\mathbf{i})}_{Z,k}}{\sigma_{Z,k}^{2}}\right)-\ln{\left(\frac{|{\widehat{\Sigma}}^{(\mathbf{i})}_{\mathbf{Z}}|}{\prod_{k}\sigma_{Z,k}^{2}}\right)}-n\right]\\ +\frac{1}{2}\mathbb{E}_{\mathbf{i}\sim p(\mathbf{i}),z\sim\hat{P}^{(\mathbf{i})}\left(Z\right)}\left[\frac{\left(f(z)-\hat{\mu}^{(\mathbf{i})}_{Y}-\left(\widehat{\bm{c}}^{(\mathbf{i})}_{\mathbf{Z}Y}\right)^{\top}\left(\widehat{\Sigma}^{(\mathbf{i})}_{\mathbf{Z}}\right)^{-1}(\mathbf{z}-\hat{\bm{\mu}}^{(\mathbf{i})}_{Z})\right)^{2}}{\sigma_{Y|Z}^{2}}\right.\\ \left.+\frac{{\widehat{\sigma^{2}}}^{(\mathbf{i})}_{Y}-\left(\widehat{\bm{c}}^{(\mathbf{i})}_{\mathbf{Z}Y}\right)^{\top}\left(\widehat{\Sigma}^{(\mathbf{i})}_{\mathbf{Z}}\right)^{-1}\widehat{\bm{c}}^{(\mathbf{i})}_{\mathbf{Z}Y}}{\sigma_{Y|Z}^{2}}-\ln{\left(\frac{{\widehat{\sigma^{2}}}^{(\mathbf{i})}_{Y}-\left(\widehat{\bm{c}}^{(\mathbf{i})}_{\mathbf{Z}Y}\right)^{\top}\left(\widehat{\Sigma}^{(\mathbf{i})}_{\mathbf{Z}}\right)^{-1}\widehat{\bm{c}}^{(\mathbf{i})}_{\mathbf{Z}Y}}{\sigma_{Y|Z}^{2}}\right)}-1\right]\,.$ (29)
The overall algorithm is described in the Algorithm 1 of main text.
@@ Appendix F EXPERIMENTAL DETAILS
F.1 Experimental settings
@@
For the adjacency matrix, we sample all non-zero entries uniformly in the interval $[-1,1]$. For general adjacency matrices, the lower triangular elements of the adjacency matrix are non-zero, where we assume that the target $Y$ has only incoming edges and the variables are arranged in topological order. For chain graphs, the first lower off-diagonal entries are non-zero. The exogenous variables $\mathbf{U}$ and shift interventions $\mathbf{i}$ are independent Gaussian with $U_{j},i_{j}\sim\mathcal{N}(0,1)$ for $j=1,...,n$.
Double well
@@
We solve the system of ODEs numerically on a grid of 101 time points $t_{k}$ for $k=0,\dots,100$ equally spaced between $t=0$ and $t=40$ using a numerical integration method. The initial conditions are $x(0)=-2.07414285+5\times 10^{-7}\times\varepsilon_{x}$, with $\varepsilon_{x}\sim\mathrm{Uniform}(-1,1)$ and $v(0)=11$. The initial values are chosen such that there is a non-zero chance that the ball ends up in the left or right well without any additional interventions.
@@
F.2 Additional results
Figure 6: Comparison of consistency loss during training for chain graphs with and without homogeneity regularization. The training consistency loss, averaged over 20 runs, is shown for two regularization settings. Similarly, we depict the normalized ${\bm{\omega}}_{1}$ parameters. We compare two training settings: with and without homogeneity regularization (9) ($\eta_{h}=0$ or $\eta_{h}=1$, respectively). Each run corresponds to a draw of the adjacency parameters.
Linear chain
As hinted at in Section 5.1, adding homogeneity regularization (9), while promoting interpretability, comes at the cost of a worse consistency of the high-level abstraction, which is shown in Figure 6.
Double well
Figure 7: Double well experiment with homogeneity regularisation ($\eta_{h}=1$). Comparison of samples and inferred causal factors in the double well experiment. (a) Depicts the experimental setup with a ball moving in a double well potential subject to linear friction. (b) Displays the pushforward density of the high-level cause for the two settings: one where no intervention is applied (unintervened), and the other with an applied shift intervention. (c) and (d) represent the learned parameters, $\tau$ and $\omega$, respectively. (e) Illustrates the conditional mean $\mathbb{E}[Y\mid Z]$ for the inferred high-level causal mechanism. (f) and (g) show samples from the unintervened setting and the corresponding estimated density. (h) Depicts the estimated density for the intervened setting.
As shown in Figure 7, introducing homogeneity regularization (9) aligns the $\bm{\tau}$- and ${\bm{\omega}}$ maps, while still capturing the correct system dynamics. In Figure 8, during training the total loss and its consistency and homogeneity contributions converge.
Figure 8: Double well training ($\eta_{h}=1$).
Two branch linear graph
Figure 9: Comparison of learned $\bm{\tau}$- and ${\bm{\omega}}$ parameters for the two branch graph and a two-cause TCR. The two-branch, low-level graph, the variable groups for the group sparsity regularization and high-level causal model are shown on the left. The normalized $\bm{\tau}_{i}$ parameters, shown on the right, ($(\bm{\tau}_{i})_{k}/\sum_{k^{\prime}}(\bm{\tau}_{i})_{k^{\prime}}$, where $k$ corresponds to the low-level variable index for $i=1,2$) are averaged over 20 runs. Each run corresponds to a draw of the adjacency parameters. Similarly, we depict the normalized ${\bm{\omega}}$ parameters.
In order to test the linear TCR algorithm 1 for multiple high-level causes, we sample from a linear Gaussian low-level model with a graph with two branches $X_{1}\to X_{2}\to X_{3}\to X_{4}\to X_{10}\leftarrow X_{9}\leftarrow\ldots\leftarrow X_{5}$. Here, $X_{10}$ is the target variable. We train a two-cause TCR and use the two branches as groups for the group sparsity regularization (18): $G_{1}=\{1,\ldots,4\}$ and $G_{2}=\{5,\ldots,9\}$. The trained $\bm{\tau}$- and ${\bm{\omega}}$ parameters are shown in Figure 9. We observe that the regularization, together with the prior information on the relevant variable groups, correctly guides TCR to identify the two branches as the high-level causes of the target $X_{10}$, where $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$ summarize the first branch $X_{1}\to\ldots\to X_{4}$, whereas $\bm{\tau}_{2}$ and ${\bm{\omega}}_{2}$ focus on the branch $X_{5}\to\ldots\to X_{9}$.
F.3 Training parameters and computing resources
For TCR training, we use the ADAM optimizer and train on NVIDIA RTX-6000 GPUs on our internal cluster. The other training parameters are summarized in Table 1. The parameters were chosen manually to ensure the training loss converges.
Parameters Linear Double Well
learning rate $\lambda$ $0.001$ $0.001$
@@ F.3 Training parameters and computing resources
training epochs $N_{\mathrm{ite}}$ $100$ $100$-$400$
@@
intervention batch size $B_{i}$ $64$ $64$
Table 1: Experimental parameters and settings

### v2

@@ E.1 Gaussian consistency loss
$\mathrm{KL}(P||Q)=\frac{1}{2}\left[(\mu_{Q}-\mu_{P})^{\top}\Sigma_{Q}^{-1}(\mu_{Q}-\mu_{P})+\mbox{tr}(\Sigma_{Q}^{-1}\Sigma_{P})-\log\frac{|\Sigma_{P}|}{|\Sigma_{Q}|}-n\right]\,.$
@@
$\mathcal{L}_{\mathrm{cons}}=\mathbb{E}_{\mathbf{i}\sim p(\mathbf{i})}\left[\mathrm{KL}_{z}(\hat{P}^{(\mathbf{i})}(z)|\hat{P^{(\omega(\mathbf{i}))}}(z))\right]+\mathbb{E}_{z\sim\hat{P}^{(\mathbf{i})}\left(\mathbf{Z}\right)}\left[\mathrm{KL}_{Y}\left(\hat{P}^{(\mathbf{i})}\left({Y}|{\mathbf{Z}}=z\right)||{P}^{(0)}\left({Y}|{\mathbf{Z}}=z\right)\right)\right]\\ =\frac{1}{2}\mathbb{E}_{\mathbf{i}\sim p(\mathbf{i})}\left[\sum_{k}\left(\frac{({\mu}_{Z,k}+{\bm{\omega}}_{k}^{\top}\mathbf{i}-\hat{\mu}^{(\mathbf{i})}_{Z,k})^{2}}{\sigma_{Z,k}^{2}}+\frac{{\widehat{\sigma^{2}}}^{(\mathbf{i})}_{Z,k}}{\sigma_{Z,k}^{2}}\right)-\ln{\left(\frac{|{\widehat{\Sigma}}^{(\mathbf{i})}_{\mathbf{Z}}|}{\prod_{k}\sigma_{Z,k}^{2}}\right)}-n\right]\\ +\frac{1}{2}\mathbb{E}_{\mathbf{i}\sim p(\mathbf{i}),z\sim\hat{P}^{(\mathbf{i})}\left(Z\right)}\left[\frac{\left(f(z)-\hat{\mu}^{(\mathbf{i})}_{Y}-\left(\widehat{\bm{c}}^{(\mathbf{i})}_{\mathbf{Z}Y}\right)^{\top}\left(\widehat{\Sigma}^{(\mathbf{i})}_{\mathbf{Z}}\right)^{-1}(\mathbf{z}-\hat{\bm{\mu}}^{(\mathbf{i})}_{Z})\right)^{2}}{\sigma_{Y|Z}^{2}}\right.\\ \left.+\frac{{\widehat{\sigma^{2}}}^{(\mathbf{i})}_{Y}-\left(\widehat{\bm{c}}^{(\mathbf{i})}_{\mathbf{Z}Y}\right)^{\top}\left(\widehat{\Sigma}^{(\mathbf{i})}_{\mathbf{Z}}\right)^{-1}\widehat{\bm{c}}^{(\mathbf{i})}_{\mathbf{Z}Y}}{\sigma_{Y|Z}^{2}}-\ln{\left(\frac{{\widehat{\sigma^{2}}}^{(\mathbf{i})}_{Y}-\left(\widehat{\bm{c}}^{(\mathbf{i})}_{\mathbf{Z}Y}\right)^{\top}\left(\widehat{\Sigma}^{(\mathbf{i})}_{\mathbf{Z}}\right)^{-1}\widehat{\bm{c}}^{(\mathbf{i})}_{\mathbf{Z}Y}}{\sigma_{Y|Z}^{2}}\right)}-1\right]\,.$ (38)
@@ Appendix F Experimental Details
F.1 Linear Experiments
Parameters Linear (Fig. 2(a)) Two Branch (Fig. 2(b))
learning rate $\lambda$ $10^{-3}$ $10^{-3}$
learning rate scheduler - cosine annealing
No. repeated train. runs per seed $1$ $10$
simulation paths $n_{\mathrm{sim}}$ $10,000$ $10,000$
training epochs $N_{\mathrm{ite}}$ $100$ $600$
simulation batch size $B$ $128$ $128$
intervention batch size $B_{i}$ $64$ $512$
overlap reg. $\eta_{\mathrm{ov}}$ (12) 00 $0.1$
balancing reg. $\eta_{\mathrm{bal}}$ (13) 00 $10^{-3}$
Table 1: Experimental parameters and settings for the linear Gaussian experiments.
@@ F.1 Linear Experiments
For the adjacency matrix, we sample all non-zero entries uniformly in the interval $[-1,1]$. For general adjacency matrices, the lower triangular elements of the adjacency matrix are non-zero, where we assume that the target $Y$ has only incoming edges and the variables are arranged in topological order. For the two-branch graph, values in the adjacency are set to zero accordingly. For chain graphs, the first lower off-diagonal entries are non-zero. The exogenous variables $\mathbf{U}$ and shift interventions $\mathbf{i}$ are independent Gaussian with $U_{j},i_{j}\sim\mathcal{N}(0,1)$ for $j=1,...,N$.
Data and Training
The data and training parameters are summarized in Table 1. All simulation data is generated before training and reused in each epoch. We split the data into training ($70\%$), validation and test ($15\%$ each). Since the training of the two-variable model would occasionally get stuck in local minima, we run each training with 10 different random initializations of the weights and select the model with the best total validation loss (14) at the end of training. Furthermore, we use a cosine annealing learning rate scheduler with a final learning rate of $10^{-5}$.
F.2 Double Well
Simulation
@@ F.2 Double Well
We solve the system of ODEs numerically on a grid of 101 time points $t_{k}$ for $k=0,\dots,100$ equally spaced between $t=0$ and $t=10$ using a numerical integration method. The initial conditions are $x(0)=-2.07414285+5\times 10^{-7}\times\varepsilon_{x}$, with $\varepsilon_{x}\sim\mathrm{Uniform}(-1,1)$ and $v(0)=11$. The initial values are chosen such that there is a non-zero chance that the ball ends up in the left or right well without any additional interventions.
@@
Data and Training
Parameters Double Well (Fig. 3)
learning rate $\lambda$ $5\cdot 10^{-4}$
learning rate scheduler -
No. repeated train. runs per seed $1$
simulation paths $n_{\mathrm{sim}}$ $10,000$
training epochs $N_{\mathrm{ite}}$ $200$
simulation batch size $B$ $128$
intervention batch size $B_{i}$ $64$
overlap reg. $\eta_{\mathrm{ov}}$ (12) 00
balancing reg. $\eta_{\mathrm{bal}}$ (13) 00
Table 2: Experimental parameters and settings for the double well experiments.
The data and training parameters are summarized in Table 1. All simulation data is generated before training and reused in each epoch. We split the data into training ($70\%$), validation and test ($15\%$ each).
F.3 Spring-Mass System
Parameters 4 masses with different weights (Fig. 4) 2 groups of masses (Fig. 7)
learning rate $\lambda$ $10^{-4}$ $10^{-3}$
learning rate scheduler cosine annealing cosine annealing
No. repeated train. runs per seed $5$ $5$
@@ F.3 Spring-Mass System
training epochs $N_{\mathrm{ite}}$ $4,800$ $1,800$
@@
intervention batch size $B_{i}$ $64$ $512$
overlap reg. $\eta_{\mathrm{ov}}$ (12) $0.1$ $0.1$
balancing reg. $\eta_{\mathrm{bal}}$ (13) $0.1$ $0.1$
spring constant $k$ $10^{-3}$ $10^{-3}$
rest length $u_{0}$ 1 1
masses $m_{i}$ $(0.5,0.83,0.17,1.5)$ all $1$
Table 3: Experimental parameters and settings for the spring mass system experiments.
Simulation
Let $M$ be the number of masses. Then, $m_{i}\in\mathbb{R}$, $\tilde{\vec{x}}_{i}(t)\in\mathbb{R}^{2}$ and $\vec{v}_{i}(t)\in\mathbb{R}^{2}$ represent the weight, position and velocity of mass $i=1,\ldots,M$ at time $t$. $A\in\{0,1\}^{M\times M}$ is the adjacency matrix encoding the spring connections, where $A_{ij}=1$ indicates that a spring connects masses $i$ and $j$. The rest length at which the springs exert no force is denoted by $u_{0}$ and $k$ is the spring constant. Both $u_{0}$ and $k$ are assumed to be the same for all springs.
The total force acting on mass $i$ at time $t$ is given by
$\vec{F}_{i}(t)=-k\sum_{j,A_{ij}=1}\left(\|\vec{u}_{ij}(t)\|-u_{0}\right)\frac{\vec{u}_{ij}(t)}{\|\vec{u}_{ij}(t)\|}$ (41)
where $\vec{u}_{ij}(t)=\vec{x}_{i}(t)-\vec{x}_{j}(t)$ is the displacement vector from mass $j$ to mass $i$. The equations of motion are
$\frac{\mathrm{d}\tilde{\vec{x}}_{i}(t)}{\mathrm{d}t}=\vec{v}_{i}(t),\qquad\frac{\mathrm{d}\vec{v}_{i}(t)}{\mathrm{d}t}=\vec{a}_{i}(t),\quad\text{with}\quad\vec{a}_{i}(t)=\frac{\vec{F}_{i}(t)}{m_{i}}\,.$ (42)
We assume that the masses have no volume and do not collide or interact other than the forces coming from the springs.
We solve the system of ODEs numerically on a grid of 21 time points $t_{k}$ for $k=0,\dots,20$ equally spaced between $t=0$ and $t=100$ using a numerical integration method. The positions are initially set on a grid to $\tilde{\vec{x}}_{1}(t=0)=(0,0)+\tilde{\vec{x}}_{\mathrm{offset}}$, $\tilde{\vec{x}}_{2}(t=0)=(1,0)+\tilde{\vec{x}}_{\mathrm{offset}}$, $\tilde{\vec{x}}_{3}(t=0)=(0,1)+\tilde{\vec{x}}_{\mathrm{offset}}$ and $\tilde{\vec{x}}_{4}(t=0)=(1,1)+\tilde{\vec{x}}_{\mathrm{offset}}$, where $\tilde{\vec{x}}_{\mathrm{offset}}\sim\mathcal{N}(0,10)$ is a random offset that shifts the entire system. The initial velocities are independently drawn as $\vec{v}_{i}(t=0)\sim\mathcal{N}(0,0.01)$. We apply random independent velocity shifts $\Delta\vec{v}_{i}(t_{k})\sim\mathcal{N}(0,0.005)$ at each time step and integrate it into the ODE solver in the same way as for the double well experiment in App. F.2.
The feature vectors $\mathbf{X}$ used to learn the TCR of the spring-mass system consists of all velocity values for all masses across all simulated time points. The interventions $\mathbf{i}$ are the corresponding velocity interventions.
Data and Training
The data and training parameters are summarized in Table 3. All simulation data is generated before training and reused in each epoch. We split the data into training ($70\%$), validation and test ($15\%$ each). Similar to the experiments on the two-branch linear graph in App. F.1, we repeat the training runs with different weight initializations and use a cosine annealing learning rate scheduler.
F.4 Additional results
F.4.1 Spring-Mass System without Regularization
Figure 6: Spring-mass system experiment without regularization. Same experimental setup as described in Sec. 5.3 and App. F.3 with the regularization turned off, i.e. $\eta_{\mathrm{ov}}=\eta_{\mathrm{bal}}=0$. The learned high-level mechanism is $f(\mathbf{Z})\approx-0.180Z_{1}+0.125Z_{2}$.
When running the TCR algorithm without regularization, it cannot be ensured that the found solutions correspond to different properties of the low-level system, as shown in Fig. 6. There is significant mixing among the high-level variables, in particular the velocity in $x$-direction of the masses towards the end of the simulation appears in both high-level variables.
F.4.2 Grouped Spring-Mass System
Figure 7: Spring-mass system experiment with two groups of masses. (a) Simulated system of eight point masses with equal weights connected by springs in two groups of 4 and with random initial velocity (blue arrows). In contrast to the experiment shown in Sec. 5.3, target of the simulation is the center of mass speed in $x$-direction. (b-e) Learned $\bm{\tau}$- and ${\bm{\omega}}$-weights corresponding to velocity components in $x$- and $y$-direction for a TCR with two high-level variables. The learned high-level mechanism is $f(\mathbf{Z})\approx-0.0866Z_{1}-0.0782Z_{2}$.
We simulate two groups of four masses as shown in Fig. 7(a). In contrast to the experiment shown in Sec. 5.3, all masses have equal weight and the target is the center of mass velocity in $x$-direction at the end of the simulation. The data and training parameters are summarized in Table 3.
Since the only interactions between masses are mediated by the springs, as described in App. F.3, the two groups of masses do not influence each other and are thus fully independent. The learned TCR identifies the two groups of masses as the two independent causes of the target. This is reflected in the parameters shown in Fig. 7 (b-e), where high-level variable $Z_{1}$ is predominantly influenced by the behavior of the second group (yellow) and variable $Z_{2}$ by the first group (blue). Furthermore, we observe that the $y$-component of the velocity, which is irrelevant for the target here, is ignored by the TCR and filtered out.